# Qualitative Study for LLM-assisted Design Study Process: Strategies, Challenges, and Roles

### v2

@@ 8 Analysis: Reflect
The “Reflect” stage emphasizes critical evaluation of research progress, tools, and methodologies to identify limitations and improve outcomes. Advanced researchers used LLMs to simulate the role of a reviewer, challenging their research projects based on the abstract and identifying potential limitations simulator. While this approach provided an initial critique, participants noted that the feedback was often too generic, focusing on broad issues such as unclear motivation, scalability, or generalizability of the proposed design. To address this, participants adopted two practices to make the feedback more specific and actionable. First, they fed the LLM their own published papers to help the model learn the typical way of limitations relevant to their research domain. Second, they provided the LLM with a more specific direction, such as asking it to consider potential solutions for model fine-tuning.
Frequent LLM users highlighted the difficulty of capturing user experiences verbally via the evaluation process, particularly when investigating the usability of tools. To address this, they used LLMs to analyze performance data and assess whether the tool was suitable for real-world users assistant programmer. For example, participants prompted LLMs to generate code that could track cursor events within a visualization system, such as the duration of cursor hover in a view, click frequency of the SVG elements, or movement trajectories, which can be used to evaluate the usability of each view, with longer hover times and higher repetitive clicking frequencies generally indicating lower usability.
P13 specifically mentioned using LLMs to generate reflection summaries based on user interview scripts assistant. Specifically, researchers can ask LLM to provide the key insights and provide with a concise summary of user feedback by analyzing the transcripts.
@@ 9 Analysis: Write
The “Write” stage focuses on using LLMs to assist in drafting academic papers, reports, and summaries. Novice students often leverage LLMs to generate papers based on high-level summaries they provided for each sentence, typically written in their native (non-English) language for international students assistant. While this approach lowered the language barrier, it introduced several challenges. First, the generated text often lacked detail, insight, or unique understanding, as the LLM tended to directly translate the student’s input. Second, when using their native language to create summaries, LLM frequently struggled to capture the key points of a sentence, resulting in the misunderstanding of the user’s intention. Participants at intermediate, proficient, and expert levels took a more structured approach to writing. More specifically, they provided high-level summaries for entire sections, allowing the LLM to draft longer passages assistant. These drafts were then manually fine-tuned to ensure accuracy and alignment with the overall paper. A common challenge across all expertise levels was the lack of contextual awareness for the whole paper, which led to inconsistent terminology and style compared to the rest of the paper. To address these issues, participants used a practice of feeding LLMs domain-specific papers to learn the appropriate writing style and refine the generated text for consistency accordingly.
Some frequent LLM users among visualization researchers pointed out the process of combining the two steps above into one assistant. Specifically, they first used reasoning-capable LLMs, such as DeepSeek or GPT-o1, to first generate a detailed outline for a section based on a high-level summary. These tools then expanded the outline into more detailed summaries iteratively, down to the level of individual sentences. Finally, advanced LLMs, such as GPT-4o, were used to craft each sentence based on the detailed summaries above. This multi-step process allowed for greater precision and coherence in the final output.
However, some participants, such as P26, expressed reservations about relying on LLMs for academic writing. They argued that academic writing is a generative task, requiring the creation of original content (The leap from nothing to something) based on human-distilled knowledge—an area where LLMs are less effective. Instead, they viewed LLMs as being better suited for tasks like summarizing or reporting, where the model works with existing material to produce concise summaries. This perspective highlights that novice writers may benefit most from LLM-generated drafts, because they tend to use LLMs for tasks that align with the model’s strengths, such as summarization.
@@ 5 Guideline Summarization
Assistant: As the assistant, LLMs can aid researchers in filling gaps in domain-specific terminology and foundational concepts (Learn) and support collaboration by evaluating visualization techniques or generating recruitment emails for collaborators (Winnow). They can help summarize complex workflows, extract design requirements (Discover). Additionally, LLMs can assist in translating domain-specific knowledge into actionable design insights, decomposing requirements into sub-tasks, and providing design inspiration (Design). LLMs can analyze performance bottlenecks, and summarize user feedback to evaluate system usability (Deploy). Finally, they can summarize user interviews, and assist in drafting and refining academic papers (Reflect, Write).
@@ 6.1.1 Toward a Project-Oriented LLM Copilot
Current LLMs provide fragmented and localized assistance in design study projects, lacking the ability to capture the entire workflow holistically. As discussed in Section 4, many of the challenges associated with LLM-assisted strategies stem from their inability to maintain an understanding of the project’s global context. For example, LLMs lose track of prior information when starting a new session and require users to repeatedly reintroduce the same content. Additionally, the narrative-driven prompts are often imprecise to LLMs, making it difficult for the model to fully grasp the actual user intention. If possible, by integrating with various technical stacks—such as Overleaf for writing, VS Code for coding, or Figma for design—advanced LLMs could provide contextualized assistance through text writing, coding, or visualization design. A future direction is to develop LLM systems that can naturally integrate into the entire design study workflow.
@@ 6.1.2 LLMs for Automated Task Decomposition
Another promising direction for LLM-assisted design studies is the development of agent-based LLMs capable of autonomously decomposing course-grained tasks. Instead of requiring users to manually break down tasks or provide detailed prompts, these advanced LLMs could operate based on high-level commands, automatically decomposing them into sub-tasks and executing them seamlessly. For example, given a set of collected design requirements, an agent-based LLM could independently perform task abstraction, and then it can generate the outputs, such as code from global visual representations (e.g., views) to detailed visual encodings (e.g., glyphs).
@@
6.1.3 Are Human Researchers Indispensable in Design Study Process?
While LLMs have shown significant strength in assisting with various aspects of design studies, all participants commented that it is unlikely that they will fully replace human researchers in the foreseeable future. First, the process of brainstorming or assessing ideas cannot be entirely taken over by LLMs. Even though LLMs have a vast corpus of visualization knowledge, they struggle to determine whether an idea is actually “interesting” or not. Human researchers, on the other hand, leverage their subjective judgment, taste, and personal domain expertise to evaluate ideas based on higher-level abstractions built upon their accumulated knowledge in visualization.
Second, while LLMs excel at summarizing existing knowledge, they fall short in creative thinking. Although there have been advancements in enabling LLMs to generate creative outputs, the majority of their results remain generic and lack the originality required for groundbreaking insights. Tasks that need critical thinking or deep reflection are areas where LLMs are less capable, while human researchers can provide profound and unique perspectives that are important to advancing the field.
Finally, experts agreed that one of the most challenging parts of visualization design for LLMs is interaction. The success or failure of an interaction is highly dependent on the subjective experiences of domain users, which vary significantly across different systems and tasks. Evaluating and refining interactions requires iterative collaboration between visualization researchers and real users, thus it remains an open challenge for LLMs.
@@ 6.2.1 Trust and Usability Concerns
Although most participants in our study expressed a positive attitude towards LLMs and their potential to support design studies, there are also concerns regarding trust and usability regarding LLMs. For example, some participants, such as P21, highlighted the “dark side” of LLM usage, where users may think of these tools as unreliable for specific tasks. Future work should focus on understanding the reasons behind these negative perceptions and exploring how LLMs can be improved to address these concerns. Second, although we have sought to include participants from diverse regions and research backgrounds, these participants may already have a predisposition toward using LLMs that do not represent the whole population of design study researchers. Third, while the paper proposes guidelines for LLM usage in design studies, it does not empirically validate these recommendations. Future work should involve testing these guidelines in real-world scenarios.
6.2.2 Rapid Evolution of LLMs
This paper focuses on the current generation of LLMs and how they are being used in design studies. However, it is obvious that LLMs are evolving rapidly, with continuous improvements in their capabilities and applications. Thus, some of the insights presented in this work may become outdated as newer and more advanced LLMs emerge. Despite this limitation, we have provided deep discussions and practical guidelines for the design study field. For example, we provided potential research directions for better integrating LLMs into design studies in the future. Additionally, we have discussed why humans cannot be replaced by LLMs in the foreseeable future. These contributions provide value for future studies to build upon, even as LLM technology becomes increasingly advanced.
@@ 6.2.3 Limited Focus on Domain- and LLM-Specific Usage
This paper investigated the strategies, challenges, and roles associated with integrating Large Language Models (LLMs) into the visualization-specific design study process. Through a multi-stage qualitative study involving 30 participants across various expertise levels and domains, we identified four primary roles for LLMs: Connector, Simulator, Programmer, and Assistant. Each role supports researchers differently across the nine stages of the design study methodology. Also, we also conducted post-study analysis to reveal how researchers treat LLM-assisted design study regarding the importance and difficulty of using LLMs. Moreover, we proposed the useful guidelines to benefits all domain researchers with different expertise, followed by the discussion of several implications for future research.

### v4

@@ 8 Analysis: Reflect
The “Reflect” stage emphasizes the critical analysis of the design study’s outcomes, and refining existing guidelines based on new insights. Advanced researchers used LLMs to simulate the role of a reviewer, challenging their research projects based on the abstract and identifying potential limitations simulator. While this approach provided an initial critique, participants noted that the feedback was often too generic, focusing on broad issues such as unclear motivation, scalability, or generalizability of the proposed design. To address this, participants adopted two practices to make the feedback more specific and actionable. First, they fed the LLM their own published papers to help the model learn the typical way of limitations relevant to their research domain. Second, they provided the LLM with a more specific direction, such as asking it to consider potential solutions in terms of the effectiveness of the approach used for resolving the problem.
Frequent LLM users highlighted the difficulty of capturing user experiences verbally via the evaluation process, particularly when investigating the usability of tools. To address this, they used LLMs to analyze performance data and assess whether the tool was suitable for real-world users assistant programmer. For example, participants prompted LLMs to generate code that could track cursor events within a visualization system, such as the duration of cursor hover in a view, click frequency of the SVG elements, or movement trajectories, which can be used to evaluate the usability of each view, with longer hover times and higher repetitive clicking frequencies generally indicating lower usability. P13 specifically mentioned using LLMs to generate reflection summaries based on user interview scripts assistant. Specifically, researchers can ask LLM to provide the key insights and provide with a concise summary of user feedback by analyzing the transcripts.
@@ 9 Analysis: Write
The “Write” stage focuses on documenting the findings in a coherent paper, and emphasizing clear abstractions and contributions. Novice students often leverage LLMs to generate papers based on high-level summaries they provided for each sentence, typically written in their native (non-English) language for international students assistant. While this approach lowered the language barrier, it introduced several challenges. First, the generated text often lacked detail, insight, or unique understanding, as the LLM tended to directly translate the student’s input. Second, when using their native language to create summaries, LLM frequently struggled to capture the key points of a sentence, resulting in the misunderstanding of the user’s intention. Participants at intermediate, proficient, and expert levels took a more structured approach to writing. More specifically, they provided high-level summaries for entire sections, allowing the LLM to draft longer passages assistant. These drafts were then manually fine-tuned to ensure accuracy and alignment with the overall paper. A common challenge across all expertise levels was the lack of contextual awareness for the whole paper, which led to inconsistent terminology and style compared to the rest of the paper. To address these issues, participants used a practice of feeding LLMs domain-specific papers to learn the appropriate writing style beforehand and refine the generated text for consistency accordingly.
Some frequent LLM users among visualization researchers pointed out the process of combining the two steps above into one assistant. Specifically, they first used reasoning-capable LLMs, such as GPT-o1 or DeepSeek, to first generate a detailed outline for a section based on a high-level summary. These tools then expanded the outline into more detailed summaries iteratively, down to the level of individual sentences. Finally, advanced LLMs, such as GPT-4o, were used to craft each sentence based on the detailed summaries above. This multi-step process allowed for greater coherence in the final output.
However, some participants, such as P26, expressed reservations about relying on LLMs for academic writing. They argued that academic writing is a generative task, requiring the creation of original content (like the leap from nothing to something) based on human-distilled knowledge—an area where LLMs are less effective. Instead, they viewed LLMs as being better suited for tasks like summarizing or reporting, where the model works with existing material to produce concise summaries. This perspective highlights that novice writers may benefit most from LLM-generated drafts, because they tend to use LLMs for tasks that align with the model’s strengths, such as summarization.
@@ 5 Guideline Summarization
Assistant: As the assistant, LLMs can aid researchers in filling gaps in domain-specific terminology and foundational concepts (Learn) and support collaboration by evaluating visualization techniques or generating recruitment emails for collaborators (Winnow). They can help summarize complex workflows, extract design requirements (Discover). Additionally, LLMs can assist in translating domain-specific knowledge into actionable design insights, decomposing requirements into sub-tasks, and providing design inspiration (Design). LLMs can analyze performance bottlenecks and summarize user feedback to evaluate system usability (Deploy). Finally, they can summarize user interviews, and assist in drafting and refining academic papers (Reflect, Write).
@@ 6.1.1 Toward a Project-Oriented LLM Copilot
Current LLMs provide fragmented and localized assistance in design study projects, lacking the ability to capture the entire workflow holistically. As discussed in Section 4, many of the challenges associated with LLM-assisted strategies stem from their inability to maintain an understanding of the project’s global context. For example, LLMs lose track of prior information when starting a new session and require users to repeatedly reintroduce the same content. Additionally, the narrative-driven prompts are often imprecise, making it difficult for the model to fully grasp the actual user intention. If possible, by integrating with various technical stacks—such as Overleaf for writing, Visual Studio Code for coding, or Figma for design—advanced LLMs could provide contextualized assistance through text writing, coding, or visualization design accordingly. Hence, a future direction will probably focus on developing LLM systems that can naturally integrate into the entire design study workflow.
@@ 6.1.2 LLMs for Automated Task Decomposition
Another promising direction for LLM-assisted design studies is the development of agent-based LLMs capable of autonomously decomposing coarse-grained tasks. Instead of requiring users to manually break down tasks or provide detailed prompts, these advanced LLMs could operate based on high-level commands, automatically decomposing them into sub-tasks and executing them seamlessly. For example, given a set of collected design requirements, an agent-based LLM could independently perform task abstraction, and then it can generate the outputs, such as code from global visual representations (e.g., views) to detailed visual encodings (e.g., glyphs) in each view.
@@
6.1.3 Are Human Researchers Indispensable in the Design Study Process?
While LLMs have shown significant strength in assisting with various aspects of design studies, all participants commented that it is unlikely that they will fully replace human researchers in the foreseeable future. First, the process of brainstorming or assessing ideas cannot be entirely taken over by LLMs. Even though LLMs have a vast corpus of visualization knowledge, they struggle to determine the quality of an idea, for example, telling whether it is actually “interesting” or not. Human researchers, on the other hand, leverage their subjective judgment, taste, and personal domain expertise to evaluate ideas based on higher-level abstractions built upon their accumulated knowledge in visualization.
Second, while LLMs excel at summarizing existing knowledge, they fall short in creative thinking. Although there have been advancements in enabling LLMs to generate creative outputs, the majority of their results remain generic and lack the originality required for groundbreaking insights. Tasks that need critical thinking or deep reflection are areas where LLMs are less capable, while human researchers can provide profound and insightful perspectives that are important to advancing the field.
Third, experts agreed that one of the most challenging parts of visualization design for LLMs is interaction. The success or failure of an interaction is highly dependent on the subjective experiences of domain users, which vary significantly across different systems and tasks. Evaluating and refining interactions requires iterative collaboration between visualization researchers and real users, thus it remains an open challenge for LLMs.
@@ 6.2.1 Trust and Usability Concerns
Although most participants in our study expressed a positive attitude towards LLMs and their potential to support design studies, there are also concerns regarding trust and usability regarding LLMs. For example, some participants, such as P21, highlighted the “dark side” of LLM usage, where users may think of these tools as unreliable for specific tasks, such as fine-tuning the mining models integrated in an interactive system. Second, although we have sought to include participants from diverse regions and research backgrounds, these participants may already have a predisposition toward using LLMs that do not represent the whole population of design study researchers. Third, while the paper proposes guidelines for LLM usage in design studies, it does not empirically validate these recommendations. Future work should involve assessing these guidelines in real-world scenarios.
6.2.2 Dynamic Progression of LLMs
This paper focuses on the current generation of LLMs and how they are being used in design studies. However, it is obvious that LLMs are evolving rapidly, with continuous improvements in their capabilities and applications. Thus, some of the insights presented in this work may become outdated as more advanced LLMs emerge. Despite this limitation, we have provided deep discussions and practical guidelines for the design study field. For example, we provided potential research directions for better integrating LLMs into design studies in the future. Additionally, we have discussed why humans cannot be replaced by LLMs in the foreseeable future. These contributions provide value for future studies to build upon, even as LLM technology becomes increasingly advanced.
@@ 6.2.3 Limited Focus on Domain- and LLM-Specific Usage
This paper investigated the strategies, challenges, and roles associated with integrating Large Language Models (LLMs) into the visualization-specific design study process. Through a multi-stage qualitative study involving 30 participants across various expertise levels and domains, we identified four primary roles for LLMs: Connector, Simulator, Programmer, and Assistant. Each role supports researchers differently across the nine stages of the design study methodology. Also, we revealed the attitudes towards how researchers treat LLM-assisted design studies regarding the importance and difficulty of using LLMs through the post-study analysis. Moreover, we proposed specific useful guidelines to benefit all domain researchers with different expertise, followed by a discussion of the potential implications for future research.